OpenStudy (anonymous):

\[\frac{ 2+ \sqrt{2} }{ \sqrt{2 - 1} }\]

OpenStudy (anonymous):

\[\frac{ 2 + \sqrt{2} }{ \sqrt{2 - 1} }\]

OpenStudy (anonymous):